Trains A and B are traveling in the same direction on parallel
track. Train A is traveling at 80 miles per hour and train B is
traveling at 100 miles per hour. Train A passes a station at 3:20 am.
If train B passes the same station at 3:35 am ., at what time will
train B catch up to train A?

Answer by mananth(16949)

the distance traveled by both is same when one catches up with the other.
Let them meet at distance x
Train A speed= 80 mph , starts at 03:20 pm
Train B speed= 100 mph starts at 03:35 pm

Time difference between the starting of the two= 0.25 hours

Time taken by Train A = X / 80
Time taken by Train B = X / 100

X / 80 - X / 100 = 0.25

LCD - 400
Multiply equation by 400
400 / 80 x- 400 / 100 x= 0.25 * 400
5 x - 4 x = 100
1 x = 100
/ 1
x= 100 miles
Train A speed= 80 mph
Distance = 100 miles
Time taken = 100 / 80
Time taken = 1.25 hour

They will be together at 04:35 pm
Train B catches up at 4:35 pm

m.ananth@hotmail.ca